OpenStudy (el_arrow):

test the series for convergence or divergence. Use the comparison or limit comparison test

OpenStudy (el_arrow):

\[\sum_{k=1}^{\infty} \frac{ 3\sqrt{k}+2 }{ \sqrt{k^3 + 3k^2 +1}}\]

OpenStudy (el_arrow):

help

OpenStudy (perl):

you can try limit comparison

OpenStudy (el_arrow):

would it be an = the equation above and bn= \[\frac{ 3k^\frac{ 1 }{ 2 }}{ k^\frac{ 3 }{ 2 } }\]

OpenStudy (el_arrow):

?

OpenStudy (perl):

yes I would compare it to sum 1/ k ^(3/2)

OpenStudy (perl):

a direct comparison also works

OpenStudy (displayerror):

Can you also simplify that last expression to \(\frac{3}{k}\) and compare to the series \(\frac{1}{k}\)?

OpenStudy (perl):

$$ \frac{ 3\sqrt{k}+2 }{ \sqrt{k^3 + 3k^2 +1}} \leq \frac{3\sqrt k + 2}{\sqrt{k^3}} $$

OpenStudy (perl):

and the right side can be split apart

OpenStudy (anonymous):

i like the eyeball test

OpenStudy (el_arrow):

whats the eyeball test

OpenStudy (anonymous):

the degree of the denominator needs to be larger than the degree of the numerator by more than one

OpenStudy (el_arrow):

is that why he has the +2 there?

OpenStudy (el_arrow):

for the bn not the an

OpenStudy (anonymous):

the degree of the numerator is \[\frac{1}{2}\] and the degree of the denominator is \[\frac{3}{2}\] so no way

OpenStudy (el_arrow):

so what happens if the degree of the numerator is less than the denominator

OpenStudy (perl):

$$ \large { \frac{ 3\sqrt{k}+2 }{ \sqrt{k^3 + 3k^2 +1}} \leq \frac{3\sqrt k + 2}{\sqrt{k^3}} = \frac{3}{\sqrt {k^2}} + \frac{2}{\sqrt{k^3}} \\ = \frac{3}{k} + \frac{2}{k^{3/2}} } $$ Now we see that the direct comparison test fails , because the right side diverges as a series. (harmonic series). We can use limit comparison test. and by limit comparison test it diverges.

OpenStudy (perl):

it is easier if we simplify the original series

OpenStudy (el_arrow):

why are diving by \[\sqrt{k}\]

OpenStudy (perl):

or we can try satellites solution. subtract the degree of the denominator minus the degree of the numerator.

OpenStudy (el_arrow):

cant it says to do it using comparison or limit comparison

OpenStudy (perl):

if you want to use the limit comparison test. this is it: $$ \Large { \lim_{k \to \infty} \frac{\frac{ 3\sqrt{k}+2 }{ \sqrt{k^3 + 3k^2 +1}} }{\frac1k} } $$

OpenStudy (el_arrow):

okay and i cancel out the 1/k making it \[\frac{ 3+2 }{ \sqrt{k^2+3k+1} }\]

OpenStudy (el_arrow):

right?

OpenStudy (perl):

if you want to use the limit comparison test. this is it: $$ \Large { \lim_{k \to \infty} \frac{\frac{ 3\sqrt{k}+2 }{ \sqrt{k^3 + 3k^2 +1}} }{\frac1k} \\ \therefore \\ \lim_{k \to \infty} {\frac{ 3\sqrt{k}+2 }{ \sqrt{k^3 + 3k^2 +1}} }\cdot \frac k 1 } $$

OpenStudy (el_arrow):

so the numerator is 3k+2k = 5k right

OpenStudy (perl):

i dont see how how you got that, the step before either

OpenStudy (el_arrow):

multiply the k times the sqrt(k) equals k and 3*k is 3k

OpenStudy (el_arrow):

no my bad its not that

OpenStudy (perl):

$$ \Large { \lim_{k \to \infty} \frac{\frac{ 3\sqrt{k}+2 }{ \sqrt{k^3 + 3k^2 +1}} }{\frac1k} \\ \therefore \\ \lim_{k \to \infty} {\frac{( 3\sqrt{k}+2) }{ \sqrt{k^3 + 3k^2 +1}} }\cdot \frac k 1 \\ \lim_{k \to \infty} {\frac{ 3\sqrt{k}\cdot k +2k }{ \sqrt{k^3 + 3k^2 +1}} } } $$

OpenStudy (el_arrow):

its \[3k \sqrt{k}+2k\]

OpenStudy (perl):

right :)

OpenStudy (el_arrow):

so after that do i just plug in infinity for k

OpenStudy (perl):

that will give you infinity / infinity

OpenStudy (perl):

but we can divide top and bottom by sqrt(k^3)

OpenStudy (el_arrow):

really?

OpenStudy (perl):

$$ \Large { \lim_{k \to \infty} \frac{\frac{ 3\sqrt{k}+2 }{ \sqrt{k^3 + 3k^2 +1}} }{\frac1k} \\ \therefore \\ \lim_{k \to \infty} {\frac{( 3\sqrt{k}+2) }{ \sqrt{k^3 + 3k^2 +1}} }\cdot \frac k 1 \\ \lim_{k \to \infty} {\frac{ 3\sqrt{k}\cdot k +2k }{ \sqrt{k^3 + 3k^2 +1}} } \\ \therefore \\ \lim_{k \to \infty} {\frac{ 3\sqrt{k}\cdot k +2k }{ \sqrt{k^3 + 3k^2 +1}} }\cdot \frac{\frac{1}{\sqrt{k^3}}}{\frac{1}{\sqrt{k^3}}} } $$

OpenStudy (perl):

also note that : \( \Large 3\sqrt{k}\cdot k = 3~ k^{1/2}\cdot k = 3 k^{3/2} = 3 \sqrt{k^3} \)

OpenStudy (el_arrow):

oh yes i see

OpenStudy (el_arrow):

your dividing everything by the largest degree right?

OpenStudy (perl):

correct

OpenStudy (perl):

$$ \Large { \lim_{k \to \infty} \frac{\frac{ 3\sqrt{k}+2 }{ \sqrt{k^3 + 3k^2 +1}} }{\frac1k} \\ \therefore \\ \lim_{k \to \infty} {\frac{( 3\sqrt{k}+2) }{ \sqrt{k^3 + 3k^2 +1}} }\cdot \frac k 1 \\ \lim_{k \to \infty} {\frac{ 3\sqrt{k}\cdot k +2k }{ \sqrt{k^3 + 3k^2 +1}} } \\ \therefore \\ \lim_{k \to \infty} {\frac{ 3\sqrt{\color{red}{k^3}} +2k }{ \sqrt{k^3 + 3k^2 +1}} }\cdot \frac{\frac{1}{\sqrt{k^3}}}{\frac{1}{\sqrt{k^3}}} \\ \therefore \\ \lim_{k \to \infty} {\frac{ 3\sqrt{ \frac{k^3}{k^3}} +2\cdot \frac{k}{\sqrt{k^3}} }{ \sqrt{\frac{k^3}{k^3} + 3\frac{k^2}{k^3}+\frac{1}{k^3}}} } \\ \therefore \\ \lim_{k \to \infty} {\frac{ 3\cdot 1 +2\cdot \frac{1}{\sqrt{k}} }{ \sqrt{1 + \frac{3}{k}+\frac{1}{k^3}}} } } $$

OpenStudy (perl):

now you can plug in infinity

OpenStudy (el_arrow):

it will equal to 1 right

OpenStudy (perl):

not quite

OpenStudy (perl):

$$ \Large { \lim_{k \to \infty} \frac{\frac{ 3\sqrt{k}+2 }{ \sqrt{k^3 + 3k^2 +1}} }{\frac1k} \\ \iff \\ \lim_{k \to \infty} {\frac{( 3\sqrt{k}+2) }{ \sqrt{k^3 + 3k^2 +1}} }\cdot \frac k 1 \\ \\ \iff \\ \lim_{k \to \infty} {\frac{ 3\sqrt{k}\cdot k +2k }{ \sqrt{k^3 + 3k^2 +1}} } \\ \iff \\ \lim_{k \to \infty} {\frac{ 3\sqrt{\color{red}{k^3}} +2k }{ \sqrt{k^3 + 3k^2 +1}} }\cdot \frac{\frac{1}{\sqrt{k^3}}}{\frac{1}{\sqrt{k^3}}} \\ \iff \\ \lim_{k \to \infty} {\frac{ 3\sqrt{ \frac{k^3}{k^3}} +2\cdot \frac{k}{\sqrt{k^3}} }{ \sqrt{\frac{k^3}{k^3} + 3\frac{k^2}{k^3}+\frac{1}{k^3}}} } \\ \iff\\ \lim_{k \to \infty} {\frac{ 3\cdot 1 +2\cdot \frac{1}{\sqrt{k}} }{ \sqrt{1 + \frac{3}{k}+\frac{1}{k^3}}} } = {\frac{ 3\cdot 1 +2\cdot \frac{1}{\sqrt{\infty }} }{ \sqrt{1 + \frac{3}{\infty}+\frac{1}{\infty^3}}} } \\ = {\frac{ 3\cdot 1 +2\cdot \color{red}0 }{ \sqrt{1 + \color{red}0+ \color{red}0}} } = \color{blue}3 } $$

OpenStudy (el_arrow):

so it converges since its greater than 1 right?

OpenStudy (perl):

by the limit comparison test, both of the series have the same behavior. if one diverges, the other diverges. if one converges, the other converges. and we know that $$ \Large \sum_{k=1}^{\infty} \frac{1}{k} $$ diverges

OpenStudy (perl):

therefore the other series must diverge as well .

OpenStudy (el_arrow):

how did we get 1/k again?

OpenStudy (perl):

that part you had to use reasoning or 'eyeballing'

OpenStudy (perl):

because overall the degree of the numerator was 1/2. the degree of the denominator was 3/2 and 3/2 - 1/2 = 2/2 = 1 . so you will have 1\k

OpenStudy (el_arrow):

oh okay

OpenStudy (perl):

$$ \Large \frac {k^{1/2}} {k^{3/2}} = \frac{1}{k^{3/2 - 1/2}} = \frac{1}{k^{2/2}} = \frac1k $$

OpenStudy (el_arrow):

is it the same when we have \[\frac{ 1+3^n }{ 1+4^n} = \frac{ 3^n }{ 4^n }\]

OpenStudy (perl):

$$\Large \frac{ 1+3^n }{ 1+4^n} \sim \frac{ 3^n }{ 4^n } = \left( \frac{3}{4}\right)^n \text{, for large n} $$

OpenStudy (el_arrow):

or \[\frac{ 1+n^2 }{ 2n+1 } = \frac{ n^2 }{ 2n } = \frac{ n }{ 2}\]

OpenStudy (perl):

yes , they are equivalent for large n . and that clearly diverges

OpenStudy (perl):

$$ \Large{ \frac{ 1+n^2 }{ 2n+1 } \sim \frac{ n^2 }{ 2n } = \frac{ n }{ 2} } $$

OpenStudy (el_arrow):

n/2 would diverge as well correct?

OpenStudy (perl):

you can show it rigorously using limits and dividing as we did above by the highest degree

OpenStudy (perl):

correct, $$\Large{ \lim_{n \to \infty} \frac n2 = \frac{\infty}{2}= \infty } $$

OpenStudy (el_arrow):

you mean i can divide the equation by 1/n like we did last problem

OpenStudy (perl):

so that series has no chance of converging, since the limit of the a_n diverges.

OpenStudy (perl):

correct

OpenStudy (el_arrow):

okay thanks a lot perl

OpenStudy (perl):

yw :)

OpenStudy (el_arrow):

how long are you going to be on for cause i might have questions later

OpenStudy (perl):

I will be here for another hour or so. Also you can post and tag me and i will come to it later , if i'm not here.

OpenStudy (el_arrow):

okay thanks